Solve. 10x – 0.62 = 9(x – 0.95)
Answer by fractalier(6550)

From
10x – 0.62 = 9(x – 0.95)
we expand the right side and collect like terms, then solve...
10x - 0.62 = 9x - 8.55
x = -7.93